calculate the PH of 40ml 0.2M CH3COOH and 40 ml 0.2M NaOH 

Grade:

1 Answers

Kaushal Negi
4 years ago
The pH will be 7
As
M = \frac{M1V1 - M2V2}{V1 + V2} 
M = \frac{0.2*40 - 0.2*40}{40 + 40}
M = \frac{0 - 0}{80}
M = 0
So it is a neutral solution and for neutral solution, pH is 7
Please approve